OpenStudy (anonymous):

how do I divide this polynomial? 8y^6-3^2-4y over 4y please show work

OpenStudy (anonymous):

(8y^6-3^2-4y)/(4y) = (8y^6)/(4y) + (-3^2)/(4y) + (-4y)/(4y)

OpenStudy (anonymous):

then what?

OpenStudy (anonymous):

y^a/y^b = y^(a-b) (8y^6)/(4y) = 8/4 * y^6/y = 2*y^(6-1) = 2*y^5

OpenStudy (anonymous):

I don't get it

OpenStudy (the_fizicx99):

http://www.wolframalpha.com/input/?i=8y%5E6-3%5E2-4y+%2F+4y Look at the step by step. Sorry I'm doing homework myself. :/

OpenStudy (anonymous):

1. y^a/y^b = y^(a-b) 2. (8y^6)/(4y) = 8/4 * y^6/y 3. y^6/y=y^5 4. 8/4 * y^6/y = 2*y^5 which step is confusing?

OpenStudy (anonymous):

just the way youre typing it

OpenStudy (anonymous):

I don't get it